OpenStudy (artemis17):

The shape of a roller coaster is modeled by a polynomial function, R(x). Describe how to find the x-intercepts of R(x) and how to construct a rough graph of R(x) so that the engineer can predict when there will be no change in the direction of the coaster. You may create a sample polynomial to be used in your explanations. Gets a medal!!!!

OpenStudy (xguardians):

1. You can find the x-intercepts of R(x) by setting it equal to zero and solving for x. 2. You can prevent the coaster from changing direction by using a simple polynomial that is always positive (positive slope).

OpenStudy (artemis17):

What do I set equal to zero though?

OpenStudy (xguardians):

If you have the equation of a line as y = 5x + 3 Set y = 0 0 = 5x + 3 -3 = 5x x = -3/5 At that point the y-value is 0.

OpenStudy (artemis17):

So to find the x-intercepts of R(x), I can set y to zero like you did?

OpenStudy (xguardians):

Yep, exactly.

OpenStudy (artemis17):

Then all I have to do is write how to graph r(x) which basically x=-3/5 and the y value?

OpenStudy (xguardians):

So take the previous function I had, y = 5x + 3 Then plug in the x-value we found by setting y = 0 y = 5(-3/5) + 3 y = -15/5 + 3 y = -3 + 3 y = 0 Therefore, you get zero again.

OpenStudy (artemis17):

ohh okay!! Thank you helps a lot..